# Global well-posedness of non-integrable hyperbolic-ellptic Ishimori system in the critical Sobolev space

### v1

@@ Abstract.
We consider the Cauchy problem for the hyperbolic-elliptic Ishimori system with general decoupling constant $\kappa\in\mathbb{R}$ and prove global well-posedness in the critical Sobolev space. The proof relies primarily on new bilinear estimates, which are established via a novel div-curl lemma first introduced by the second author in [14]. Our approach combines the caloric gauge technique with $U^{p}$-$V^{p}$ type Strichartz estimates to handle the hyperbolic structure of the equation. The results extend previous work on the integrable case $\kappa=1$ to general $\kappa$ and provide a unified framework that also applies to hyperbolic and elliptic Schrödinger maps in dimensions $d\geq 2$.
@@ 1. Introduction
The system exhibits different behaviors depending on the coupling constant. When $\kappa=0$, the system reduces to the two-dimensional hyperbolic Heisenberg equation (hyperbolic Schrödinger map). When $\kappa=1$, the system is completely integrable and can be solved via the inverse scattering method. For general $\kappa\in\mathbb{R}$, the system is non-integrable and presents additional analytical challenges.
The local and global regularity properties of the Cauchy problem associated with the hyperbolic-elliptic Ishimori system have been extensively studied, see [6, 9, 13]. Soyeur [9] established local and global existence for small initial data in $H^{3}$, and uniqueness for large data in $H^{4}$. Wang [13] proved local well-posedness for small data in $H^{\sigma}_{Q}$ for $\sigma>\frac{3}{2}$. For the integrable case $\kappa=1$, Bejenaru, Ionescu, and Kenig [6] established global well-posedness for small data in the critical Sobolev space $\dot{H}^{1}_{Q}(\mathbb{R}^{2})$. For a more comprehensive historical overview, we refer to [8, Section 9.2].
Regarding the elliptic Schrödinger map equation, which relates to the special case $\kappa=0$ in (1.1), substantial progress has been made on low-regularity local and global well-posedness, see [1, 3, 4]. Of particular relevance to our approach is the seminal work of Bejenaru, Ionescu, Kenig, and Tataru [1], who proved global existence for the elliptic Schrödinger map in critical Sobolev spaces for dimensions $d\geq 2$. Their proof relies crucially on the caloric gauge formulation and local smoothing estimates for the associated linear flow.
This paper aims to extend these results to the non-integrable Ishimori system ($\kappa\in\mathbb{R}$) using the caloric gauge framework from [1]. However, there are significant differences: the local smoothing estimates used in the elliptic case are not directly applicable due to the hyperbolic structure of the Ishimori equation. Instead, we employ $U^{p}$-$V^{p}$ type Strichartz estimates and a new div-curl lemma first introduced by the second author in [14]. This novel div-curl lemma is crucial for establishing the necessary bilinear estimates, particularly for controlling high-low and low-high frequency interactions in Bony’s paraproduct decomposition. We note that our methods are also applicable to both hyperbolic and elliptic Schrödinger maps in dimensions $d\geq 2$, providing a unified approach.
@@
which is equipped with the metric $d_{H^{\sigma}_{Q}}(f,g)=\|f-g\|_{H^{\sigma}}$. Similarly for $\dot{H}^{\sigma}_{Q}$ and $H^{\sigma,\rho}_{Q}$. We define
@@ 1.1. The modified spin model in caloric gauge
Expanding the covariant derivatives yields the following nonlinear ultrahyperbolic Schrödinger equation
@@
where $\mu_{l}=\delta_{1l}-\delta_{2l}$ reflects the ultrahyperbolic signature.
@@ 1.2. Outline of the proof
([1, Prop. 4.2])Given arbitary interval $I\subseteq\mathbb{R}$ and $S\in C(I;H^{\infty}_{Q})$ satisfying the smallness condition
@@
(1.50) $a_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}b_{k}(\sigma),\ \mathcal{D}(P_{k}\bm{\psi})\leq 2^{-\sigma k}b_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
@@
(1.51) $a_{k}(\sigma)$ $\lesssim b_{k}(\sigma),\ \sigma\in[0,\sigma_{1}]$
@@
The second bootstrap proposition concerns the ultrahyperbolic Schrödinger evolution.
Proposition 1.7 (Ultrahyperbolic Schrödinger bootstrap assumptions).
@@
(1.53) $\mathcal{D}(P_{k}\bm{\psi})\leq b_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
@@
(1.54) $b_{k}(\sigma)+\mathcal{D}(P_{k}\bm{\psi})\lesssim c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
@@
([11, Prop. 3.13]) Given $S_{0}^{h}\in H^{\infty}_{Q}$ with $\|S^{h}_{0}\|_{\dot{H}^{1}}\ll 1$ for $h=0,1$. There exists exists a smooth one-parameter family $S^{h}_{0}\in C^{\infty}_{h}([0,1];H^{\infty}_{Q})$ such that
@@ 2.1. Multilinear expression
then for any bilinear form $L$ with integrable kernel we have
@@ 2.2. Linear estimates
Let $G=L^{4}_{t,x}\cap L^{\infty}_{t}L^{2}_{x}$ denote the standard Strichartz space. To establish Strichartz estimates for solutions to the linear ultrahyperbolic Schrödinger equation
@@
The spaces $U^{p}_{\text{UH}}$ and $V^{p}_{\text{UH}}$ are adaptations of the standard $U^{p}$ and $V^{p}$ spaces to the specific dispersive properties of the ultrahyperbolic Schrödinger operator $i\partial_{t}+\partial_{1}^{2}-\partial_{2}^{2}$. Their construction follows the general theory developed in [7, Chap. 4], where analogous spaces are introduced for various dispersive equations. The key properties—embeddings, duality, and connection to Strichartz estimates—are proved using the methods outlined therein.
@@ 2.3. Bilinear estimates
$\quad\quad=\Re(\partial_{m}\overline{v_{2}}\Xi_{m}(D)P_{k_{2}}N^{\prime})-\frac{1}{2}\partial_{m}\Re(\overline{v_{2}}\Xi_{m}(D)P_{k_{2}}N^{\prime})=2^{k_{2}}L(P_{k_{2}}v,P_{k_{2}}N^{\prime}).$
@@
$\sup_{y}\iint\Big(\int\Re(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}\int\Re(\overline{v_{2}}\partial_{m}v_{2})-\int\Im(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}\int\Im(\overline{v_{2}}\partial_{m}v_{2})\mathrm{d}\widehat{x}_{m}\Big)\mathrm{d}t\mathrm{d}x_{m}$
@@ 3. Estimates in the heat flow direction
$\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))\lesssim\sum_{k^{\prime}\leq k}\langle 2^{2k^{\prime}+2j}\rangle^{-n}\beta_{k^{\prime},j}$
@@
(3.62) $\|P_{k}[(\bm{A}(s)\cdot\bm{A}(s),\ \bm{\psi}(s)\cdot\bm{\psi}(s),\ \partial_{x}\bm{A}(s))\cdot\psi_{t}(s)]\|\lesssim B_{2}\varepsilon 2^{k}2^{-2j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}.$
@@
(3.68) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
@@
Denote by $k_{\min},k_{\text{med}}$ and $k_{\max}$ the minimum, median and maximum of the tuple $(k_{1},k_{2},k)$ respectively. For summation over $|k_{\text{med}}-k_{\max}|\leq 4$ and $|k_{2}-k|\geq 100$, we have
@@
For summmation over $|k_{2}-k|\leq 100$, we have $|k_{1}-k^{\prime}|\leq 200$ and thus
@@
$\lesssim\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}2^{\frac{k}{2}+k^{\prime}}2^{-\sigma k^{\prime}}\mathcal{D}(P_{k^{\prime}}w)b_{k^{\prime}}b_{k^{\prime}}(\sigma)(2^{2k^{\prime}}r)^{-\frac{1}{8}}\langle 2^{2k^{\prime}}r\rangle^{-6}\mathrm{d}r$
$\lesssim\langle 2^{2k}s\rangle^{-3}\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{3-N}2^{\frac{k}{2}+k^{\prime}}2^{-\sigma k^{\prime}}\mathcal{D}(P_{k^{\prime}}w)b_{k^{\prime}}b_{k^{\prime}}(\sigma)(2^{2k^{\prime}}r)^{-\frac{1}{8}}\langle 2^{2k^{\prime}}r\rangle^{-3}\mathrm{d}r$
@@
$\lesssim\mathcal{D}(P_{k^{\prime}}w)2^{-\sigma k}2^{-\frac{k^{\prime}-k}{2}}\langle 2^{2k}s\rangle^{-3}b_{k}b_{k}(\sigma),$
@@
(3.76) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
(3.77) $\sup_{y}\|P_{k}\partial_{x}\bm{\psi}(s)P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-4}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi}),$
@@
(3.78) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi}).$
@@
(3.79) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum_{k_{1},k_{2},m}\big\|\|P_{k}(P_{k_{1}}\bm{A}(s)P_{k_{2}}\bm{\psi}(s))\|_{L^{\infty}_{\hat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{\hat{x}_{m}}}\big\|_{L^{2}_{t,x_{m}}}$
$\lesssim\sum_{k_{1},k_{2},m}2^{\frac{1}{2}(\min\{k,k_{1},k_{2}\}+k)}\|P_{k_{1}}\bm{A}(s)\|_{L^{\infty}_{t,x_{m}}L^{2}_{\hat{x}_{m}}}\big\|\|P_{k_{2}}\bm{\psi}(s)\|_{L^{2}_{\hat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{\hat{x}_{m}}}\big\|_{L^{2}_{t,x_{m}}}$
$\lesssim 2^{-\frac{k^{\prime}}{2}}\sum_{k_{1},k_{2}}2^{\frac{1}{2}(\min\{k,k_{1},k_{2}\}+k)}2^{\frac{k_{1}}{2}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi})$
@@
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-\sigma k}b_{k}(\sigma)2^{\frac{3k}{4}}s^{-\frac{1}{8}}\langle 2^{2k}s\rangle^{-\frac{13}{4}}\mathcal{D}(P_{k^{\prime}}\bm{\psi})$
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-(\sigma-1)k}b_{k}(\sigma)(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}\bm{\psi}),$
@@
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-(\sigma-1)k}b_{k}(\sigma)(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k_{3}}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}\bm{\psi}).$
@@
(3.81) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{1},k_{2}\}}\|P_{k_{1}}\bm{A}(s)\|_{G}\|P_{k_{2}}\bm{\psi}(s)\|_{G}\|P_{k^{\prime}}\bm{\psi}\|_{G}$
$\lesssim\mathcal{D}(P_{k^{\prime}}\bm{\psi})\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{1},k_{2}\}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)(2^{2k_{1}}s)^{-\frac{1}{8}}\langle 2^{2k_{1}}s\rangle^{-\frac{13}{4}}\langle 2^{2k_{2}}s\rangle^{-4}$
$\lesssim 2^{-\sigma k}2^{\frac{k^{\prime}}{2}}s^{-\frac{1}{4}}\mathcal{D}(P_{k^{\prime}}\bm{\psi})\sum_{k_{1},k_{2}}2^{\frac{1}{8}(\min\{k,k_{1},k_{2}\}-\max\{k,k_{1},k_{2}\})}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
@@
$\lesssim 2^{-(\sigma-1)k}2^{\frac{k^{\prime}-k}{2}}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi}),$
where we use the fact that $\max\{k_{1},k_{2}\}\geq k$. This gives the stated bound when $k^{\prime}\leq k$ and completes the proof.

### v2

# arXiv:2601.03576v2[math.AP] 01 Mar 2026

@@ Abstract.
We consider the Cauchy problem for the hyperbolic-elliptic Ishimori system with general decoupling constant $\kappa\in\mathbb{R}$ and prove global well-posedness in the critical Sobolev space. The proof relies primarily on new bilinear estimates, which are established via a novel div-curl lemma first introduced by the second author in [13]. Our approach combines the caloric gauge technique with $U^{p}$-$V^{p}$ type Strichartz estimates to handle the hyperbolic structure of the equation. The results extend previous work on the integrable case $(\kappa=1)$ to general $\kappa$ and provide a unified framework which also works for the hyperbolic and elliptic Schrödinger maps in dimensions $d\geq 2$.
@@ 1. Introduction
Assuming that $S(t,x)$ is a solution to the Ishimori system (1.1) with initial data $S_{0}$, then the scaled function $S_{\lambda}(t,x)=S(\lambda^{2}t,\lambda x)$ is also a solution with initial data $S_{0}(\lambda x)$. A direct scaling computation gives
(1.3) $\|S_{\lambda}(t)\|_{\dot{H}^{s}(\mathbb{R}^{d})}=\lambda^{s-\frac{d}{2}}\|S(t)\|_{\dot{H}^{s}(\mathbb{R}^{d})},$
so that in the two-dimensional case, the critical Sobolev space for $S$ is $\dot{H}^{1}(\mathbb{R}^{2})$.
The local and global regularity properties of the Cauchy problem associated with the hyperbolic-elliptic Ishimori system have been extensively studied, see [5, 8, 12]. Soyeur [8] established local and global existence for small initial data in $H^{3}$, and uniqueness for large data in $H^{4}$. Wang [12] proved local well-posedness for small data in $H^{\sigma}_{Q}$ (see (1.4) below) for $\sigma>\frac{3}{2}$. A broader historical overview may be found in [7, Sect. 9.2].
Two special cases are of particular interest. When $\kappa=1$, the system is completely integrable. In this setting, Bejenaru, Ionescu, and Kenig [1] established global well-posedness for small data in the critical Sobolev space $\dot{H}^{1}_{Q}(\mathbb{R}^{2})$. Their analysis hinges on the observation that in the Coulomb gauge the magnetic terms vanish and the nonlinearities become non-derivative. When $\kappa=0$, the system reduces to the two-dimensional hyperbolic Heisenberg equation, equivalent to the hyperbolic Schrödinger map. Of particular relevance to our approach is the seminal work of Bejenaru, Ionescu, Kenig, and Tataru [2], who proved global existence for the elliptic Schrödinger map in critical Sobolev spaces for dimensions $d\geq 2$. Their proof relies crucially on the caloric gauge formulation and local smoothing estimates for the associated linear flow.
In this paper, we extend these results to the non-integrable Ishimori system with general $\kappa\in\mathbb{R}$ using the caloric gauge framework of [2]. However, this is highly non-trivial. Unlike the integrable case $\kappa=1$, the system for general $\kappa$ is non-integrable and its nonlinearities are derivative-dependent, introducing significant analytic obstacles. Moreover, compared to the elliptic Schrödinger map, the hyperbolic nature of the Ishimori equation prevents direct use of local smoothing estimates, which were essential in [2]. Instead, we employ Strichartz estimates in $U^{p}$-$V^{p}$ space (first introduced by Tataru in an unpublished work. See [6] for detailed introduction), and a new div-curl lemma first introduced by the second author in [13]. This new div-curl lemma plays a crucial role in establishing bilinear estimates, especially for controlling high-low and low-high frequency interactions in Bony’s paraproduct decomposition. High-high frequency interactions are handled by the $U^{p}$-$V^{p}$ type Strichartz estimates. We note that our methods also work for both ultrahyperbolic and elliptic Schrödinger maps in dimensions $d\geq 2$, providing a unified approach.
@@
which is equipped with the metric $d_{H^{\sigma}_{Q}}(f,g)=\|f-g\|_{H^{\sigma}}$. Similarly for $\dot{H}^{\sigma}_{Q}$. We define
@@ 1.1. The modified spin model in caloric gauge
Expanding the covariant derivatives yields the following nonlinear hyperbolic Schrödinger equation
@@
where $\mu_{l}=\delta_{1l}-\delta_{2l}$ reflects the hyperbolic signature.
@@ 1.2. Outline of the proof
([2, Prop. 4.2])Given arbitary interval $I\subseteq\mathbb{R}$ and $S\in C(I;H^{\infty}_{Q})$ satisfying the smallness condition
@@
(1.51) $a_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}b_{k}(\sigma),\ \mathcal{D}(P_{k}\bm{\psi})\leq 2^{-\sigma k}b_{k}(\sigma),\ \text{for}\ \sigma\in[0,\sigma_{1}],$
(1.52) $\|P_{k}\psi_{t}\|_{G}\lesssim 2^{k}\varepsilon^{\frac{1}{2}}.$
@@
(1.53) $a_{k}(\sigma)$ $\lesssim b_{k}(\sigma),\ \text{for}\ \sigma\in[0,\sigma_{1}]$
(1.54) $\|\psi_{t}(s)\|_{G}$ $\lesssim\langle 2^{2k}s\rangle^{-4}\varepsilon^{\frac{1}{2}},$
@@
The second bootstrap proposition concerns the hyperbolic Schrödinger evolution.
Proposition 1.7 (Hyperbolic Schrödinger bootstrap assumptions).
@@
(1.56) $\mathcal{D}(P_{k}\bm{\psi})\leq b_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}c_{k}(\sigma),\ \text{for}\ \sigma\in[0,\sigma_{1}],$
(1.57) $\|P_{k}\psi_{t}\|_{G}\leq 2^{k}\varepsilon^{\frac{1}{2}}.$
@@
(1.58) $b_{k}(\sigma)+\mathcal{D}(P_{k}\bm{\psi})\lesssim c_{k}(\sigma),\ \text{for}\ \sigma\in[0,\sigma_{1}],$
(1.59) $\|P_{k}\psi_{t}\|_{G}\lesssim 2^{k}\varepsilon.$
@@
([10, Prop. 3.13]) Given $S_{0}^{h}\in H^{\infty}_{Q}$ with $\|S^{h}_{0}\|_{\dot{H}^{1}}\ll 1$ for $h=0,1$. There exists exists a smooth one-parameter family $S^{h}_{0}\in C^{\infty}_{h}([0,1];H^{\infty}_{Q})$ such that
@@ 2.1. Multilinear expression
then for any trilinear form $L$ with integrable kernel we have
@@ 2.2. Linear estimates
Let $G=L^{4}_{t,x}\cap L^{\infty}_{t}L^{2}_{x}$ denote the standard Strichartz space. To establish Strichartz estimates for solutions to the linear hyperbolic Schrödinger equation
@@
The spaces $U^{p}_{\text{UH}}$ and $V^{p}_{\text{UH}}$ are adaptations of the standard $U^{p}$ and $V^{p}$ spaces to the specific dispersive properties of the hyperbolic Schrödinger operator $i\partial_{t}+\partial_{1}^{2}-\partial_{2}^{2}$. Their construction follows the general theory developed in [6, Chap. 4], where analogous spaces are introduced for various dispersive equations. The key properties—embeddings, duality, and connection to Strichartz estimates—are proved using the methods outlined therein.
@@ 2.3. Bilinear estimates
$\quad\quad=\int\Re(\partial_{m}\overline{v_{2}}\Xi_{m}(D)P_{k_{2}}N^{\prime})-\frac{1}{2}\partial_{m}\Re(\overline{v_{2}}\Xi_{m}(D)P_{k_{2}}N^{\prime})\mathrm{d}\widehat{x}_{m}=2^{k_{2}}\int L(P_{k_{2}}v,P_{k_{2}}N^{\prime})\mathrm{d}\widehat{x}_{m}.$
@@
$+\sup_{y}\iint\Big(\int\Re(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}\int\Re(\overline{v_{2}}\partial_{m}v_{2})-\int\Im(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}\int\Im(\overline{v_{2}}\partial_{m}v_{2})\mathrm{d}\widehat{x}_{m}\Big)\mathrm{d}t\mathrm{d}x_{m}$
@@ 3. Estimates in the heat flow direction
$\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))\lesssim\sum_{k^{\prime}\leq k}\langle 2^{2k^{\prime}+2j}\rangle^{-n}2^{k^{\prime}}\beta_{k^{\prime},j}$
@@
(3.62) $\|P_{k}[(\bm{A}(s)\cdot\bm{A}(s),\ \bm{\psi}(s)\cdot\bm{\psi}(s),\ \partial_{x}\bm{A}(s))\cdot\psi_{t}(s)]\|\lesssim B_{2}\varepsilon^{\frac{1}{2}}2^{k}2^{-2j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}.$
@@
(3.68) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
@@
Denote by $k_{\min},k_{\text{med}}$ and $k_{\max}$ the minimum, median and maximum of the tuple $(k_{1},k_{2},k)$ respectively. For summation over $|k_{\text{med}}-k_{\max}|\leq 4$ and $k^{\prime}\geq k_{2}+100$, we have
@@
For summmation over $|k_{\text{med}}-k_{\max}|\leq 4$ and $k_{2}\geq k^{\prime}-100\geq k$, we have $|k_{1}-k_{2}|\leq 10$ and thus
@@
$\lesssim\sum_{k_{2}\geq k^{\prime}-100}\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}2^{\frac{k}{2}+k_{2}}2^{-\sigma k_{2}}\mathcal{D}(P_{k^{\prime}}w)b_{k_{2}}b_{k_{2}}(\sigma)(2^{2k_{2}}r)^{-\frac{1}{8}}\langle 2^{2k_{2}}r\rangle^{-6}\mathrm{d}r$
@@
$\lesssim\mathcal{D}(P_{k^{\prime}}w)2^{-\sigma k}2^{\frac{k^{\prime}-k}{2}}\langle 2^{2k}s\rangle^{-3}b_{k}b_{k}(\sigma),$
@@
(3.76) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
(3.77) $\sup_{y}\|P_{k}\partial_{x}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
@@
(3.78) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w).$
@@
(3.79) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum_{k_{1},k_{2},m}\big\|\|P_{k}(P_{k_{1}}\bm{A}(s)P_{k_{2}}\bm{\psi}(s))\|_{L^{\infty}_{\hat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w^{y}\|_{L^{2}_{\hat{x}_{m}}}\big\|_{L^{2}_{t,x_{m}}}$
$\lesssim\sum_{k_{1},k_{2},m}2^{\frac{1}{2}(\min\{k,k_{1},k_{2}\}+k)}\|P_{k_{1}}\bm{A}(s)\|_{L^{\infty}_{t,x_{m}}L^{2}_{\hat{x}_{m}}}\big\|\|P_{k_{2}}\bm{\psi}(s)\|_{L^{2}_{\hat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w^{y}\|_{L^{2}_{\hat{x}_{m}}}\big\|_{L^{2}_{t,x_{m}}}$
$\lesssim 2^{-\frac{k^{\prime}}{2}}\sum_{k_{1},k_{2}}2^{\frac{1}{2}(\min\{k,k_{1},k_{2}\}+k)}2^{\frac{k_{1}}{2}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)\mathcal{D}(P_{k^{\prime}}w)$
@@
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-\sigma k}b_{k}(\sigma)2^{\frac{3k}{4}}s^{-\frac{1}{8}}\langle 2^{2k}s\rangle^{-\frac{13}{4}}\mathcal{D}(P_{k^{\prime}}w)$
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-(\sigma-1)k}b_{k}(\sigma)(2^{2k}s)^{-\frac{1}{8}}\langle 2^{2k}s\rangle^{-\frac{13}{4}}\mathcal{D}(P_{k^{\prime}}w),$
@@
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-(\sigma-1)k}b_{k}(\sigma)(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}w).$
@@
(3.81) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{1},k_{2}\}}\|P_{k_{1}}\bm{A}(s)\|_{G}\|P_{k_{2}}\bm{\psi}(s)\|_{G}\|P_{k^{\prime}}w\|_{G}$
$\lesssim\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{1},k_{2}\}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)(2^{2k_{1}}s)^{-\frac{1}{8}}\langle 2^{2k_{1}}s\rangle^{-\frac{13}{4}}\langle 2^{2k_{2}}s\rangle^{-4}$
$\lesssim 2^{-\sigma k}2^{\frac{k^{\prime}}{2}}s^{-\frac{1}{4}}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{\frac{1}{8}(\min\{k,k_{1},k_{2}\}-\max\{k,k_{1},k_{2}\})}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
@@
$\lesssim 2^{-(\sigma-1)k}2^{\frac{k^{\prime}-k}{2}}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
where we use the fact that $\max\{k_{1},k_{2}\}\geq k-4$. This gives the stated bound when $k^{\prime}\leq k$ and completes the proof.